OpenStudy (osnad):

solve for a 10+[3/a]=[5^a]

OpenStudy (shubhamsrg):

\[10 + 3/a = 5^a\] do you mean this ?

OpenStudy (osnad):

yes

OpenStudy (lgbasallote):

or maybe you mean this \[\huge 10 + \frac 3a = 5^a\]

OpenStudy (osnad):

yes

OpenStudy (lgbasallote):

so which is it?

OpenStudy (osnad):

the last one

OpenStudy (anonymous):

3 months old o_o

OpenStudy (osnad):

yeah i really need it badly

OpenStudy (anonymous):

is a an integer in this ?

OpenStudy (anonymous):

only a calculator or an exact graph could solve this

OpenStudy (cwrw238):

let t = 5^a then ln t = ln 5^a = a ln 5 and a = ln t / ln 5 so we have 10 + 3ln 5/ ln t = t i'm not sure this helps

OpenStudy (cwrw238):

i think akash is right there the solution done graphically http://www.wolframalpha.com/input/?i=10+%2B+3%2Fa+%3D+5%5Ea

OpenStudy (anonymous):

I don't get it, how can we solve this algebraically?

OpenStudy (shubhamsrg):

newton's method of approximation? long and tedious ,but still useful..

OpenStudy (shubhamsrg):

[okay wait,, will [ ] symbolize greatest integer ?? makes it bit easier then i guess..

OpenStudy (osnad):

still waiting

OpenStudy (anonymous):

is that greates integer sign ????

OpenStudy (anonymous):

i love floor function :)

OpenStudy (anonymous):

what means [] here? Osnad ???

OpenStudy (osnad):

is just a bracket